OpenStudy (gabylovesyou):

Simplify 4 over quantity 2 plus 7 i.

OpenStudy (freckles):

multiply top and bottom by bottom's conjugate

OpenStudy (gabylovesyou):

8 + 28i

OpenStudy (gabylovesyou):

? @freckles

OpenStudy (ybarrap):

Let me get you started: $$ {4\over2}+7={4\over 2}+{2\times 7\over 2}={4+2\times 7\over 2} $$ Can you finish?

OpenStudy (ybarrap):

Is that an "i" in the end?

OpenStudy (ybarrap):

In that case, we have: $$ {4\over2}+7i=2+7i $$

OpenStudy (gabylovesyou):

\[\frac{ 4 }{ 2 + 7i }\]

OpenStudy (ybarrap):

$$ {4\over2+7i}={4\over2+7i}\times {2-7i\over2-7i}={8-28i\over4+49} $$

OpenStudy (ybarrap):

Note that we multiply by the complex conjugate of \(2-7i\) of the denominator because we want to make the denominator "real."

OpenStudy (gabylovesyou):

ty

OpenStudy (ybarrap):

You're welcome